# Pulse-induced memory-like effect in cyclotron motion?

### v1

@@ III Conclusions
In this brief article, we have shown how an effect somewhat similar to what is conventionally termed as electromagnetic memory can arise in a rather simple scenario. The age-old problem of a charged particle moving on a circular path due to a uniform magnetic field is modified by introducing a short-duration electric pulse. This causes a ‘permanent change’–the radius can increase/decrease, the centre of the circle always shifts away and the velocity as well as its magnitude can increase or decrease. We also showed that for a pair of particles moving in the same ${\bf B}$ and ${\bf E}$ fields but with different conditions at a given time, the relative separation, velocity do change. This is analogous to memory as found using a gravitational wave pulse where, a change in relative separation distinguishes between pulses ‘with memory’ or ‘without memory’. We have also mentioned (with some explicit values) how our results may be seen in a cyclotron with necessary practical adjustments.
@@
As mentioned in the Introduction, it is true that the example discussed here does not concern the usual ‘wave memory’. It is, rather, a completely non-relativistic scenario where the effect of a short duration pulse on a trajectory is explored. One may argue that this can indeed also happen in simple mechanical systems – say, when a particle is acted upon by an impulsive force. In particle mechanics though, the particle will actually have to be ‘struck’ in order to give an impulse. This is not the case when a charged particle moves in magnetic and electric fields–the interaction of the charge with the magnetic and electric field is not a ‘contact’ interaction and may be visualised as an ‘influence’ over a short duration, which leads to a change which persists. Even though we do not have a ‘wave field’ here, the ‘time dependent pulse’ can bring about a change akin to memory. Such a pulse-induced memory is not uncommon in condensed matter (magnetic) systems as is well-known and well-studied in the literature (see eg. jacs , acs among others).
@@
An obvious extension of this work is to try out different pulse shapes –delta function, triangular, Gaussian, sech-squared are the simplest examples –the last two being smooth functions. It may be difficult to do the smooth function examples analytically, but all scenarios are surely tractable numerically. An important related question is about whether there is a pulse for which we can observe a shape change in the trajectory–i.e. a circle going over to an ellipse or something else. Such a model will be useful because it is, in some sense, perhaps closer to the idea of memory as proposed and pursued in the context of gravitational wave physics ayansayan .

### v2

@@ IV Conclusions
In this brief article, we have shown how an effect somewhat similar to what is conventionally termed as electromagnetic memory, can arise in a rather simple scenario. The age-old problem of a charged particle moving on a circular path due to a uniform magnetic field is modified by introducing a short-duration electric pulse. This causes a ‘permanent change’–the radius can increase/decrease, the centre of the circle always shifts away and the velocity as well as its magnitude can increase or decrease. We also showed that for a pair of particles moving in the same ${\bf B}$ and ${\bf E}$ fields but with different conditions at a given time, the relative separation, velocity do change. The origins of this ‘permanent change’, as explained in the previous section, can be traced to a net difference in the vector potential which, in turn, is linked to a gauge transformation. This feature is analogous to memory as found using a gravitational wave pulse where, a change in relative separation has its origins (for asymptotically flat spacetimes) in diffeomorphisms of flat spacetime at asymptotic infinity. One may also distinguish between pulses ‘with memory’ or ‘without memory’ for the given system, using the change in relative separation as a signature. Finally, we have also suggested (with some explicit values) how our results may be seen in a cyclotron with necessary practical adjustments.
@@
As stated in the Introduction, it is true that the example discussed here does not concern the usual ‘wave memory’ caused by radiation at null infinity. It is, rather, a completely non-relativistic scenario where the effect of a short duration pulse on a trajectory is explored. One may argue that this can indeed also happen in simple mechanical systems – say, when a particle is acted upon by an impulsive force. In particle mechanics though, the particle will actually have to be ‘struck’ in order to give an impulse. This is not the case when a charged particle moves in magnetic and electric fields–the interaction of the charge with the magnetic and electric field is not a ‘contact’ interaction and may be visualised as an ‘influence’ over a short duration, which leads to a change which persists. Even though here we do not have a ‘radiative wave field’ at null infinity, the ‘time dependent pulse’ can bring about a change akin to memory. Further, we have shown how our memory-like effect is related to the asymptotic (in time) values of the vector potentials which are gauge-related. A little search in the literature reveals that such a pulse-induced memory is not uncommon in condensed matter (magnetic) systems as is well-known and well-studied in various articles (see eg. jacs , acs among others).
@@
An obvious extension of this work is to try out different pulse shapes –delta function, triangular, Gaussian, sech-squared are the simplest examples –the last two being smooth functions. It may be difficult to do the smooth function examples analytically, but all scenarios are surely tractable numerically. An important related question is about whether there is a pulse for which we can observe a shape change in the trajectory–i.e. a circle going over to an ellipse or some other closed curve. Such a model will be useful because it is, in some sense, perhaps closer to the idea of memory as proposed and pursued in the context of gravitational wave physics ayansayan . As a concluding remark we mention that it may be possible to extrapolate the basic ideas presented here to the context of gravitoelectromagnetism gem1 ; gem2 where, at a linearised level, one recovers the Maxwell and Lorentz force equations with the ‘electric’ and ‘magnetic’ fields born out of the Riemann tensor components of a background spacetime geometry.
@@ ACKNOWLEDGEMENTS
DATA AVAILABILITY STATEMENT
No Data associated in the manuscript.